### v1

# Impact of reactor antineutrinos on the neutrino floor in low-mass WIMP-like dark matter searches

Abstract
The sensitivity of conventional direct dark matter searches for weakly interacting massive particles (WIMPs) is ultimately limited by coherent elastic neutrino–nucleus scattering (CENS), which produces nuclear recoils indistinguishable from WIMP signals and defines the so-called neutrino floor. While the effects of solar, geoneutrino, diffuse supernova, and atmospheric neutrinos have been extensively studied in this context, the contribution from reactor antineutrinos has received comparatively little attention. We present the first systematic evaluation of how reactor antineutrino fluxes, modeled as a function of reactor–detector distance, modify the neutrino floor for low-mass WIMP searches using SuperCDMS-like high-voltage germanium detectors. Both discovery-limit and opacity-based formulations of the neutrino floor are examined under consistent assumptions. We find that proximity to gigawatt-scale reactors within 10 km can raise the neutrino floor by up to a few orders of magnitude, significantly reducing the sensitivity to sub-10 GeV/ dark matter. Beyond 100 km, the reactor contribution becomes negligible. These results establish reactor proximity as a critical site-dependent parameter in defining the neutrino floor and provide guidance for optimizing the design and siting of future low-threshold dark matter experiments.
I Introduction
The existence of dark matter (DM) is supported by a wide range of astrophysical and cosmological observations first_proposed_DM:1933gu; rotation_curve_1970; bullet_cluster_2006; Planck_2015_CMB. However, its particle nature remains unknown. Weakly Interacting Massive Particles (WIMPs) are theoretically well-motivated and represent one of the leading dark matter candidates WIMP_1_1984; WIMP_2_1995_supersymmetric; Relic_WIMP_abundances_2012. Direct detection experiments search for nuclear recoils produced by the elastic scattering of WIMPs off target nuclei. Since both the WIMP mass () and the WIMP–nucleon scattering cross section () are unknown, these experiments explore a wide parameter space in and CDMSliteR3_WIMP_PLR_2019; CRESST_2024_WIMP; LZ_2023_WIMP; PandaX_2025_WIMP; XENONnT_2025_WIMP.
Next-generation direct detection experiments are rapidly approaching sensitivities to very small cross sections SuperCDMS_SNOALB_projected_SNOWMASS_2021; CRESST_2025_projected_sensitivity; xenon2024xenonnt_status. At these levels, an irreducible background from coherent elastic neutrino–nucleus scattering (CENS) freedman becomes a limiting factor. In CENS, neutrinos scatter coherently off entire nuclei, producing nuclear recoils that are experimentally indistinguishable from those expected from WIMP interactions. The smallest WIMP cross section discoverable under this background corresponds to the point at which a WIMP signal would produce a statistically significant excess (typically ) over the CENS background. Below this level, the dark matter signal becomes effectively obscured, defining the so-called neutrino floor Billard:2013qya; Ruppin_followup:2014bra; Overcome_neutrino_floor:2020lva.
The neutrino floor is not a strict boundary but rather a sensitivity threshold that depends on several factors, including the theoretical understanding of CENS cross sections, the neutrino fluxes from various sources, and the total exposure of experiments. To emphasize its diffuse and site-dependent nature, the term neutrino fog has also been introduced OHare:2021utq; CJPL_neutrino_floor_Fan:2025sde. Although recent CENS measurements are consistent with Standard Model predictions COHERENT_CEvNS; COUNSplus_CEvNS; COHERENT:Ge_2025, the dominant uncertainty in determining the neutrino floor arises from systematic uncertainties in the neutrino fluxes, which can reach up to 50%.
Neutrino-induced backgrounds arise from both natural and man-made sources. Natural backgrounds include solar neutrinos, geoneutrinos, the diffuse supernova neutrino background (DSNB), and atmospheric neutrinos, while man-made contributions mainly originate from nuclear reactor operations. The flux of reactor electron antineutrinos can be substantial, particularly for experiments located near reactor facilities. Although solar and atmospheric neutrino backgrounds are well characterized in the context of the neutrino floor, the role of reactor antineutrinos has traditionally received less attention Billard:2013qya; OHare:2021utq. For searches targeting low-mass WIMPs in the sub-GeV/ to GeV/ mass range, CENS induced by reactor antineutrinos can significantly affect sensitivity, depending on the experiment’s distance from the nearest reactor.
Germanium-based cryogenic detectors operated at high bias voltages (HV) and achieving ultra-low energy thresholds, such as those employed by SuperCDMS CDMSliteR3_WIMP_PLR_2019, are particularly well suited for probing WIMPs in this mass regime. Therefore, our study focuses on determining the neutrino floor for germanium-based detectors. We systematically quantify the impact of reactor antineutrino fluxes on the neutrino floor. Using detailed spectral models of a GW reactor configuration and accounting for the detector quenching response, we evaluate how CENS-induced backgrounds vary with distance from the reactor and how they influence WIMP discovery limits using two widely adopted definitions of the neutrino floor: the discovery-based and opacity-based approaches.
The pioneering works by Billard et al. Billard:2013qya and Ruppin et al. Ruppin_followup:2014bra established the discovery-limit-based definition of the neutrino floor for xenon, silicon, and germanium targets, without accounting for reactor or geoneutrino fluxes. Subsequently, O’Hare OHare:2021utq introduced the opacity (or “neutrino fog”) framework, which included reactor backgrounds only for the Gran Sasso site, but without exploring its distance dependence. More recently, Fan et al. CJPL_neutrino_floor_Fan:2025sde evaluated the opacity-based neutrino floor at the China Jinping Underground Laboratory (CJPL), without considering distance-dependent reactor effects, while D. Aristizabal Sierra et al. reactor_anti_neutrino_contribution_PRD_2024 studied site-specific reactor antineutrino fluxes across different underground laboratories but did not assess their impact on the neutrino floor.
The present work builds upon these foundational studies by providing the quantitative assessment of how reactor fluxes—explicitly modeled as a function of reactor–detector distance—affect both discovery-limit-based and opacity-based neutrino floors for germanium detectors. This study thus establishes, for the first time, the explicit distance dependence of the neutrino floor and highlights the importance of site-specific reactor antineutrino modeling in future dark matter sensitivity projections.
Our results show that proximity to a gigawatt-scale reactor within ten kilometers can elevate the neutrino floor by up to a few orders of magnitude across the low-mass WIMP region (in the few GeV/ range), compared to baseline “no-reactor” studies Billard:2013qya; Billard:APPEC_report; Ruppin_followup:2014bra or those assuming reactor–detector separations of km OHare:2021utq. This substantial enhancement has important implications for the design and siting of next-generation direct detection experiments targeting sub-10 GeV/ dark matter.
The paper is organized as follows. Section II presents the modeling of the WIMP signal. Section III describes the modeling of CENS-induced backgrounds. In Section IV, we examine the impact of reactor antineutrinos on the neutrino floor using two widely adopted definitions. Section V discusses the implications of these effects for future dark matter experiments. Finally, Section VI summarizes the main findings and conclusions.
II WIMP Signal Rate Modeling
II.1 Differential Event Rate
Direct detection experiments primarily aim to measure the tiny nuclear recoil energy imparted to a target nucleus through the elastic scattering of a WIMP. The differential event rate per unit detector mass can be expressed as WIMP_rate_Lewin
where is Avogadro’s number, is the local dark matter density, is the atomic mass number of the target nucleus, and is the WIMP mass. The WIMP–nucleus reduced mass is , where denotes the nuclear mass. The WIMP–nucleus cross section at zero momentum transfer is represented by , and is the nuclear form factor that accounts for the loss of coherence at finite momentum transfer. The detector efficiency is assumed to be unity above the threshold.
The minimum WIMP velocity required to produce a recoil of energy is
The astrophysical dependence of the rate enters through the mean inverse speed,
where is the WIMP velocity distribution in the Galactic rest frame, is the Earth’s velocity relative to the Galactic frame, and is the Galactic escape velocity. This integral represents the fraction of WIMPs with sufficient speed to induce a recoil of energy . In the standard halo model Standard_halo_model, the Galactic WIMP velocity distribution is assumed to follow a truncated Maxwell–Boltzmann form with parameters, (escape velocity), and (Earth’s velocity). Analytic approximations follow the formalism of Lewin and Smith WIMP_rate_Lewin.
II.2 Spin-Independent WIMP–Nucleon Scattering in Germanium
We focus on studying the neutrino floor for germanium-based detectors. Naturally occurring germanium isotopes are predominantly even–even nuclei with negligible total spin, suppressing spin-dependent (SD) scattering. Consequently, the recoil spectrum is dominated by coherent spin-independent (SI) interactions, allowing SD contributions to be neglected in sensitivity projections.
The WIMP–nucleus cross section at zero momentum transfer is related to the fundamental SI WIMP–nucleon cross section by
where is the WIMP–nucleon reduced mass and is the nucleon mass. Coherent scattering off all nucleons enhances the cross section by a factor of , while the reduced-mass ratio accounts for the kinematic scaling between the nucleus and the nucleon.
The nuclear form factor is modeled using the standard Helm parameterization WIMP_rate_Lewin:
where is the spherical Bessel function of order one, is the momentum transfer, is the nuclear skin thickness, and , with and .
II.3 Quenching and the Measured WIMP Energy Spectrum
In germanium-based detectors operated in high-voltage (HV) mode, such as those used by SuperCDMS CDMSliteR3_WIMP_PLR_2019, the total measured phonon signal is dominated by Neganov–Trofimov–Luke (NTL) phonons LUKE1; LUKE2. The total energy of these NTL phonons is proportional to the amount of ionization produced by particle interactions. For nuclear recoils, only a fraction of the deposited energy is converted into ionization, while the remainder goes into the creation of primary phonons. Consequently, the nuclear recoil energy from WIMP scattering does not entirely lead to NTL phonon production.
This reduced ionization response, known as the quenching effect, reflects the lower ionization yield of nuclear recoils relative to electron recoils of the same energy. The ionization yield is commonly described by the Lindhard model lindhard_quenching_1963:
where and is the dimensionless energy parameter. The constant is material-dependent. For germanium (), recent CENS results from the CONUS+ Collaboration CONUS_reactor_anti-neutrino_observation_Ackermann:2025obx indicate that provides good agreement down to recoil energies of . We adopt this value and assume that the Lindhard model remains valid at lower recoil energies. Because both WIMP- and CENS-induced interactions produce nuclear recoils, they are expected to exhibit similar quenching behavior. Consequently, the exact choice of quenching model has little impact on the determination of the neutrino floor. However, applying quenching is important for obtaining a realistic prediction of the experimentally observable recoil spectrum.
The measured (electron-equivalent) energy is related to the recoil energy as
accounting for the quenching effect. The differential event rate in measured energy is obtained as
where the derivative , evaluated using Eq. (6), results in the conversion between the recoil and measured energy scales.
The resulting spectrum, , represents the observable WIMP-induced signal and forms the basis for sensitivity projections and comparisons with CENS-induced backgrounds.
III CENS-Induced Background Modeling
III.1 Differential Scattering Cross Section
Coherent elastic neutrino–nucleus scattering (CENS) occurs when a neutrino scatters elastically off an entire nucleus at low momentum transfer, resulting in a coherent enhancement of the cross section proportional to the square of the weak nuclear charge freedman. Within the Standard Model, the differential cross section for a neutrino of energy scattering off a nucleus of mass and producing a recoil energy can be expressed as CEvNS_Crosssection_Scholberg:2005qs:
where is the Fermi coupling constant, is the weak nuclear charge, and are the neutron and proton numbers of the target nucleus, and is the weak mixing angle. The prefactor associated with is close to zero; therefore, the CENS cross section is dominated by the neutron term, leading to a coherent enhancement approximately proportional to . The kinematic maximum recoil energy is given by
III.2 Quenching and the Measured CENS Energy Spectrum
The differential CENS event rate per unit detector mass is expressed as a sum over the differential rates from all neutrino sources:
where is the neutrino flux from source , is the minimum neutrino energy required to produce a nuclear recoil , and is the endpoint energy of the corresponding spectrum. The detector efficiency is assumed to be unity above threshold. The various neutrino sources contributing to the CENS background relevant to the neutrino floor are discussed in Sec. III.3.
In cryogenic germanium detectors, only a fraction of the nuclear recoil energy is converted into measurable phonon signals through the Neganov–Trofimov–Luke effect. This energy suppression, referred to as the quenching effect, has been described in Sec. II. The experimentally observable spectrum is obtained by folding the differential recoil rate with the quenching factor using Eq. 7, yielding the measured CENS energy distribution.
III.3 Sources of CENS Backgrounds
CENS-induced backgrounds originate from both natural and man made neutrino sources.
III.3.1 Non-Reactor Neutrino Sources
Solar Neutrinos.
Solar neutrinos are primarily produced through the chain and the CNO cycle in the Sun. They dominate the neutrino flux below 18 MeV and constitute the primary CENS background for low-mass WIMP searches ( GeV/). In this study, we include contributions from , , , 7Be, 8B, 13N, 15O, and 17F neutrinos, using the differential event rates in nuclear recoil energy for a germanium detector from Ref. Billard:2013qya.
Atmospheric Neutrinos.
Atmospheric neutrinos are produced by interactions of cosmic rays with Earth’s atmosphere and include both neutrinos and antineutrinos. These neutrinos are highly energetic, with energies extending up to the TeV scale. However, as the neutrino energy increases, the coherence of the CENS process diminishes due to the increasing momentum transfer. Following Ref. Billard:2013qya, we consider atmospheric neutrinos up to 1 GeV, which primarily obscure WIMP signals in the mass range 100 GeV/–(10 TeV/), and adopt the corresponding differential event rate in nuclear recoil energy for germanium.
Geoneutrinos.
Geoneutrinos originate from the -decays of long-lived isotopes such as uranium, thorium, and potassium within the Earth’s crust and mantle. Their flux exhibits moderate spatial variation due to differences in crustal composition and thickness. The geoneutrino flux is typically two orders of magnitude lower than the solar neutrino flux; therefore, its contribution to the CENS background is subdominant. Geoneutrinos produce low-energy recoils, primarily affecting WIMP sensitivity near GeV/. Given the moderate site-to-site variations and the overall lower flux compared to solar neutrinos, we adopt the differential CENS event rate in germanium detectors in terms of nuclear recoil energy evaluated for the China Jinping Underground Laboratory (CJPL) as a representative case Geo_neutrino_Gelmini:2018gqa.
Diffuse Supernova Neutrinos.
The diffuse supernova neutrino background (DSNB) arises from the cumulative flux of neutrinos emitted by all past core-collapse supernovae in the universe. These neutrinos have typical energies of a few MeV, with the spectrum peaking around 3–8 MeV, making them relevant to WIMP searches at intermediate masses (). We adopt the DSNB flux and the corresponding differential CENS event rate in germanium detectors as a function of nuclear recoil energy from Ref. Billard:2013qya.
All CENS recoil spectra are converted to the electron-equivalent energy scale using the same Lindhard quenching model applied to WIMP recoils. This allows a direct comparison between WIMP-induced and neutrino-induced signals in germanium detectors. The differential event rates from individual CENS sources in a germanium detector are shown in Fig. 1. As evident from the figure, the differential rate from the 8B solar neutrino component closely resembles the spectrum of a WIMP with mass 6 GeV/ and spin-independent cross section , highlighting the difficulty of distinguishing low-mass WIMP signals from neutrino-induced backgrounds.
Solar neutrinos dominate the CENS background at low WIMP masses (), while atmospheric neutrinos dominate at higher masses (). The fractional uncertainties associated with each neutrino source are summarized in Table 1. The uncertainty in the CONUS-reported quenching parameter is on the order of 1% COUNSplus_CEvNS, which is much smaller than the overall systematic uncertainties in the CENS rate. Since both CENS and WIMP-induced spectra are subject to similar quenching effects, the impact of quenching uncertainty on the neutrino floor determination is expected to be minimal and is therefore neglected in our calculations.
| Component | Uncertainty |
| pp | 0.6% |
| pep | 1% |
| hep | 30% |
| 7Be I (384 keV) | 6% |
| 7Be II (861 keV) | 6% |
| 8B | 2% |
| 13N | 15% |
| 15O | 17% |
| 17F | 20% |
| DSNB | 50% |
| Atmospheric | 25% |
| Geo (40K) | 6% |
| Geo (238U) | 30% |
| Geo (232Th) | 30% |
| Reactor | 8% |
III.3.2 Reactor Antineutrinos
Nuclear reactors are intense sources of emitted in the -decays of fission fragments from 235U, 238U, 239Pu, and 241Pu. These low-energy antineutrinos produce sub-keV recoils via CENS, overlapping the expected signal region for light WIMPs in germanium detectors.
To evaluate the reactor-induced CENS background, we consider a representative reactor similar to the Leibstadt (KKL) plant in Switzerland conus+_experiment_2024, operating at a thermal power of 3.6 GW. The reactor antineutrino flux at a detector located a distance from the reactor core is
where is the reactor thermal power, MeV is the average energy released per fission energy_per_fission, is the fission fraction of isotope , and is the emitted spectrum per fission. The flux scales as , making distance a critical parameter for near-core experiments. The fission fractions are listed in Table 2.
| Isotope | Fission Fraction |
| 235U | 53% |
| 239Pu | 32% |
| 238U | 8% |
| 241Pu | 7% |
The total reactor spectrum combines a parameterized model above 2 MeV with tabulated data below 2 MeV. For MeV, the spectra are parameterized as Hubber_reactor_spectrum_PhysRevC.84.024617; CONNIE_anti_neutrino_spectrum:2019xid
with coefficients listed in Table 3.
| Isotope | ||||
| 235U | 1.0461 | 0.870 | –0.160 | –0.0910 |
| 238U | 1.0719 | 0.976 | –0.162 | –0.0790 |
| 239Pu | 1.0527 | 0.896 | –0.239 | –0.0981 |
| 241Pu | 1.0818 | 0.793 | –0.080 | –0.1085 |
Below 2 MeV, the spectra are taken from tabulated data (Table 4):
| [MeV] | 235U | 239Pu | 238U | 241Pu |
| 0.008 | 0.024 | 0.140 | 0.089 | 0.200 |
| 0.016 | 0.092 | 0.560 | 0.350 | 0.790 |
| 0.031 | 0.350 | 2.130 | 1.320 | 3.000 |
| 0.062 | 0.610 | 0.640 | 0.650 | 0.590 |
| 0.125 | 1.980 | 1.990 | 2.020 | 1.850 |
| 0.250 | 2.160 | 2.080 | 2.180 | 2.140 |
| 0.500 | 2.660 | 2.630 | 2.910 | 2.820 |
| 0.750 | 2.660 | 2.580 | 2.960 | 2.900 |
| 1.000 | 2.410 | 2.320 | 2.750 | 2.630 |
| 1.500 | 1.690 | 1.480 | 1.970 | 1.750 |
| 2.000 | 1.260 | 1.080 | 1.500 | 1.320 |
At low energies (around 1 MeV and below), an additional contribution to the flux arises from neutron capture on 238U, with the corresponding spectrum taken from Ref. TEXONO_U238_capture:2006xds. This spectrum was normalized and scaled according to its antineutrino yield per fission, taken as 1.2 CONNIE_anti_neutrino_spectrum:2019xid. The combined reactor antineutrino spectra at different reactor–detector distances are shown in Fig. 2.
Reactor antineutrinos can constitute an important component of the CENS background for direct detection WIMP search experiments, especially for sites located near nuclear power plants. Their contribution strongly depends on the distance between the detector and the reactor core, leading to large variations in the expected event rates. As shown in Fig. 3, the CENS event rate can vary by a few orders of magnitude for detector locations within 50 km of a reactor, surpassing the solar neutrino background at distances below 10 km. Accurate modeling of the site-dependent reactor flux is therefore essential for evaluating the neutrino floor and assessing the discovery reach of future low-threshold dark matter experiments.
We validated our reactor antineutrino CENS recoil spectrum against the CENS spectrum measured by the CONUS+ experiment CONUS+_first_observation_2025. As discussed in Appendix A, our Standard Model prediction for the CENS recoil spectrum is in good agreement with the experimental data from CONUS+, measured at the same KKL reactor.
Since this study focuses on the dependence of reactor antineutrino backgrounds on distance (from a few to about 100 km), we assume that non-reactor neutrino flux components and their associated event rates remain constant when varying the detector location.
IV Impact of Reactor Antineutrinos on the Neutrino Floor
The neutrino floor defines the sensitivity threshold of direct WIMP search experiments, below which CENS backgrounds become statistically indistinguishable from potential WIMP signals. In this regime, uncertainties in the neutrino background limit the achievable discovery significance, such that further increases in exposure yield diminishing gains in sensitivity.
In this work, we examine two widely used formulations of the neutrino floor and assess the impact of reactor antineutrinos: the discovery-limit approach proposed by Billard et al. Billard:2013qya, and the opacity-based or neutrino-fog framework introduced by O’Hare OHare:2021utq.
IV.1 Discovery-Limit Based Neutrino Floor
In the discovery-limit framework Billard:2013qya, the neutrino floor is defined as the minimum WIMP–nucleon cross section for which a WIMP signal can be detected with at least significance in 90% of hypothetical experiments, in the presence of CENS backgrounds.
The detection significance is evaluated using a binned profile-likelihood ratio test, following standard methods for new-physics searches cowan2011asymptotic. Event counts in each nuclear-recoil energy bin are assumed to follow independent Poisson distributions, while uncertainties in individual neutrino flux components are treated as Gaussian nuisance parameters. The likelihood function is written as
where is the observed number of events in bin , is the predicted WIMP signal for a given dark matter mass and cross section, is the CENS background contribution from flux component , is the normalization parameter scaling the th flux component, and denotes its fractional uncertainty as provided in Tab. 1.
The corresponding conditional (background-only) and unconditional (signal-plus-background) log-likelihoods are
The profile-likelihood ratio test statistic is defined as
where are the unconditional maximum-likelihood estimates obtained by maximizing , and are the conditional maximum-likelihood estimates obtained by maximizing under the background-only hypothesis.
The discovery significance, expressed in Gaussian units, is
which quantifies the significance for rejecting the background-only hypothesis in favor of the signal-plus-background hypothesis.
To determine the discovery-limit cross section, we perform 1000 Monte Carlo pseudo-experiments for each assumed WIMP mass and cross section. Each pseudo-dataset is generated by applying independent Poisson fluctuations to the expected number of events in each energy bin, based on the combined WIMP signal and CENS background model. For each realization, the likelihood ratio and the corresponding significance are computed, and the resulting distribution captures the expected statistical variation among experiments.
The discovery-limit cross section, , is defined as the smallest cross section for which at least 90% of pseudo-experiments yield . In statistical terms, this corresponds to a power of 90% to reject the background-only hypothesis at significance, providing a robust and experimentally meaningful detection threshold.
Repeating this procedure across a range of WIMP masses yields the discovery threshold as a function of mass, and the curve connecting these thresholds defines the discovery-limit neutrino floor.
To probe a broad WIMP mass range, we perform two sets of pseudo-experiments with different detector thresholds, following the approach of Ref. Billard:2013qya. A low-threshold configuration (5.3 eVnr or 0.46 eVee using Lindhard quenching) captures the impact of solar neutrinos on low-mass WIMPs, while a high-threshold setup (7.2 keVnr or 1.65 keVee) suppresses the solar contribution and focuses on DSNB and atmospheric neutrinos relevant for high-mass WIMPs. The most sensitive discovery limit from these two regimes defines the neutrino floor across the full WIMP mass range.
The exposures used in the pseudo-experiments are chosen to provide statistically meaningful sampling of the neutrino background. For the low-threshold setup, the exposure corresponds to roughly 200 expected 8B solar neutrino events, while the high-threshold configuration yields about 500 total neutrino events Ruppin_followup:2014bra; Billard:2013qya. These correspond to exposures of approximately 0.4 ton-year and ton-year, respectively. With these choices, the statistical uncertainty on the total neutrino background remains smaller than the systematic uncertainty in the neutrino flux, allowing the pseudo-experiments to capture both statistical fluctuations and systematic uncertainties. It should be noted that these threshold and exposure values are somewhat idealized and not yet representative of current detector capabilities. Prototype detectors such as those developed by SuperCDMS HVeVR1; SuperCDMS:HVeV_compton_steps; SuperCDMS:HVeV4_DM and CRESST CRESST_2024_WIMP have achieved eV-scale thresholds but remain limited to gram-scale target masses, making ton-year exposures currently unfeasible.
All calculations incorporate the flux uncertainties summarized in Table 1 as Gaussian priors and include contributions from reactor, astrophysical, and geoneutrino CENS sources. This treatment ensures that the projected sensitivity accurately reflects the systematic uncertainties in the neutrino fluxes.
Figure 4 shows the impact of reactor antineutrinos on the neutrino floor at varying reactor–detector distances. The effect is most pronounced for dark matter masses below 10 GeV/, with the largest deviation observed around 2.5–3 GeV/. These results indicate that proximity to a reactor can substantially elevate the neutrino floor by orders of magnitude for low-mass WIMPs due to the enhanced reactor-induced CENS background. As the distance increases, the reactor contribution diminishes and the sensitivity gradually converges toward the baseline neutrino floor.
We studied how systematic uncertainties in the reactor antineutrino flux influence the neutrino floor. Figure 5 illustrates the impact of varying the systematic uncertainty in the reactor flux for two reactor–detector distances: 1 km and 10 km. For the 1 km scenario, a tenfold reduction in flux uncertainty noticeably improves the sensitivity, effectively opening up a larger region of WIMP parameter space accessible to discovery. In contrast, for a 10 km separation, the effect of flux uncertainty is much less pronounced due to the reduced reactor contribution.
Figure 6 illustrates how increasing exposure affects WIMP search sensitivity. Higher exposure enhances both the WIMP signal and background counts, reducing statistical uncertainty and improving discrimination between dark matter signals and CENS backgrounds. However, once systematic uncertainties dominate, further exposure yields diminishing returns. This limitation is most evident where the WIMP and neutrino recoil spectra are nearly degenerate—for instance, at GeV/, where the WIMP signal resembles the 8B solar neutrino spectrum. This behavior underscores the intrinsic sensitivity ceiling imposed by neutrino backgrounds.
To overcome the explicit dependence of the discovery-limit-based neutrino floor definition on exposure and threshold, a complementary definition of the neutrino floor based on the concept of opacity was introduced. This framework naturally incorporates exposure-dependent effects into the sensitivity definition.
IV.2 Opacity-Based Neutrino Floor
A more recent definition, introduced by O’Hare OHare:2021utq, characterizes the neutrino fog as the transition region where the dark matter signal spectrum becomes progressively obscured by coherent elastic neutrino–nucleus scattering backgrounds, primarily due to systematic uncertainties in neutrino fluxes. Within this framework, the neutrino floor is defined as the boundary between the Poisson-statistics–dominated regime and the systematic-uncertainty–dominated regime, quantified by the dimensionless opacity parameter.
At low exposures, discovery sensitivity improves approximately as the inverse square root of exposure, governed by Poisson fluctuations in event counts. As exposure increases, further gains become limited (Fig. 6) by systematic uncertainties in the CENS background normalization.
Following methods described in OHare:2021utq, the discovery-limit cross section at fixed detection significance , , is evaluated across a grid of WIMP masses and exposures. For each exposure, the total CENS background is obtained by summing contributions from solar, reactor, diffuse supernova, geoneutrino, and atmospheric neutrino sources. The opacity parameter, , quantifies how the discovery sensitivity degrades with increasing neutrino background (and equivalently with exposure). It is defined as the inverse logarithmic slope of the discovery-limit cross section with respect to the total CENS background OHare:2021utq:
Figure 7 (left) shows how the discovery sensitivity, expressed as , evolves with the expected number of neutrino events, revealing distinct scaling behaviors across exposure regimes. At low exposures, follows the expected dependence, where Poisson fluctuations dominate. As the exposure increases, the improvement in sensitivity gradually saturates, indicating a transition to a regime dominated by systematic uncertainties in the neutrino flux, where further gains are limited and the opacity correspondingly increases. At very high exposures, the scaling returns to the Poisson-dominated behavior, again consistent with the trend. The figure shows these features for two representative WIMP masses: 2.6 GeV/, where reactor antineutrinos form the dominant CENS background, and 6.0 GeV/, where 8B solar neutrinos majorly contribute. Differences in flux normalization and systematic uncertainties alter the position and extent of the saturation region between the two cases.
The opacity-based neutrino floor is defined as the locus in the plane where the scaling exponent , marking the transition between the statistics-limited and systematics-limited regimes. The region below this curve marks the regime where increasing exposure yields diminishing returns in sensitivity, as CENS backgrounds increasingly obscure potential WIMP signals—the so-called neutrino fog. However, since the WIMP and CENS recoil spectra are not perfectly degenerate, this limitation is not absolute; at sufficiently high exposures, the sensitivity can partially recover the Poisson-like scaling as the spectral differences become statistically distinguishable Overcome_neutrino_floor:2020lva; OHare:2021utq.
Figure 7 (right) shows how the opacity parameter evolves with exposure (top) and discovery limit (bottom) for the two WIMP masses: 2.6 GeV/ and 6.0 GeV/.
The resulting neutrino fog at varying reactor–detector distances is shown in Fig. 8, alongside the no-reactor baseline. In the absence of reactor antineutrinos, the low-mass DM sensitivity is primarily limited by solar neutrinos. An additional opaque region appears around –3 GeV/ due to the contribution from reactor antineutrinos, shifting to higher cross sections as the reactor–detector distance decreases. For distances shorter than 10 km, the reactor contribution surpasses the solar component, significantly raising the neutrino floor in the low-mass region by up to a few orders of magnitude. Unlike the discovery-limit–based floor, the opacity-based definition depends weakly on the absolute exposure choice in pseudo experiments, making the impact to the baseline floor a more robust physical effect. These results highlight that reactor antineutrino fluxes constitute a critical, site-dependent background for sub-10 GeV/ WIMP searches. Careful consideration of reactor proximity is therefore essential when designing low-threshold experiments, as even modest reactor contributions can dominate the low-energy CENS spectrum and substantially reduce dark matter discovery sensitivity.
Appendix B presents a comparison of the baseline neutrino floor—i.e., the neutrino floor obtained without including reactor antineutrinos—from this study with previously published results. The comparison is shown for both the discovery-limit–based approach Billard:APPEC_report and the opacity-based approach OHare:2021utq, focusing on Ge-based detectors. Overall, our baseline limits show reasonable agreement with previously published results under both definitions.
V Implications for Future Dark Matter Experiments
Future direct-detection experiments targeting low-mass WIMPs ( GeV/) must account for CENS backgrounds arising from nearby nuclear reactors. Reactor antineutrinos induce low-energy nuclear recoils that overlap with the expected WIMP signal region, effectively elevating the neutrino floor and reducing both sensitivity and discovery potential.
Figure 9 shows the reactor-induced modification of the neutrino floor under the discovery-limit and opacity-based definitions. Both approaches exhibit similar qualitative trends but differ by up to an order of magnitude due to their distinct conceptual treatments OHare:2021utq.
At short reactor–detector distances, the enhanced antineutrino flux dominates the low-energy recoil spectrum, substantially elevating the neutrino floor at low WIMP masses. As summarized in Table 5 for GeV/, the neutrino floor could increase by few orders of magnitude at baselines below 10 km and rapidly decreases at larger distances. We also find that the opacity-based neutrino floor is more sensitive to variations in the antineutrino flux than the discovery-limit-based definition. To minimize the impact of reactor antineutrinos (limited to within a few percent), dark matter experiments probing sub-10 GeV/ WIMPs should ideally be located more than 100 km away from a gigawatt-scale reactor. These results emphasize the importance of accurate, site-specific background modeling in future sensitivity projections.
Figure 9 also presents the current exclusion limits from leading dark matter searches, where the shaded region denotes the parameter space already ruled out experimentally. The area between the present exclusion limits and the neutrino floor corresponds to the yet unexplored parameter space accessible to conventional direct detection experiments. Future searches aim to probe this region further. As evident, reactor proximity significantly reduces this discovery window in the sub-10 GeV/ WIMP mass region, thereby constraining the potential for new discoveries in this mass range.
| Reactor–detector distance | Discovery-limit–based | Opacity–based | ||||
| [cm2] | [%] | [cm2] | [%] | |||
| 1 km | ||||||
| 5 km | ||||||
| 10 km | ||||||
| 50 km | ||||||
| 100 km | ||||||
| no reactor (baseline floor) | ||||||
Although this effect is negligible for deep underground laboratories such as SNOLAB—where typical reactor baselines exceed 100 km reactor_anti_neutrino_contribution_PRD_2024—it becomes a vital consideration for next-generation low-threshold experiments and site planning. Accurate inclusion of reactor antineutrino fluxes in sensitivity projections is therefore essential for realistic assessments of discovery reach in the sub-10 GeV/ WIMP mass regime. Mitigation strategies, including directional detection Directional_search_OHare:2015utx; Directional_search_Grothaus:2014hja, could further help preserve sensitivity in reactor-proximate environments.
VI Summary and Conclusions
We have investigated the impact of reactor antineutrinos on the sensitivity of future low-mass dark matter searches, primarily focusing on WIMPs with masses below 10 GeV/. A SuperCDMS-like high-voltage germanium detector configuration was adopted, incorporating realistic detector response via quenching corrections using Lindhard model. Two widely adopted formulations of the neutrino floor-the discovery-limit and opacity-based approaches—were evaluated under consistent assumptions to quantify the effect of reactor proximity. This study represents the first comprehensive study of the neutrino floor that explicitly incorporates site-specific reactor proximity under both statistical definitions. We find that proximity to gigawatt-scale reactors can significantly elevate the neutrino floor, particularly within 10 km, where enhancements of a few orders of magnitude arise in the sub-10 GeV/ WIMP mass range. Such elevated backgrounds from reactor antineutrinos markedly reduce the discovery reach of low-mass dark matter experiments. To suppress this effect to the few-percent level, experiments aiming to probe sub-10 GeV/ WIMPs should ideally be located at least 100 km from major reactors, beyond which the contribution becomes negligible.
These findings establish reactor proximity as an important design consideration for next-generation dark matter detectors. Future experiments exploring the low-mass WIMP parameter space must incorporate realistic, site-dependent reactor flux modeling when projecting sensitivities. Additionally, mitigation strategies such as directional detection could prove vital for maintaining sensitivity and fully exploiting the discovery potential of future low-threshold searches.
Acknowledgments
We acknowledge the support of the Department of Atomic Energy (DAE) and the Department of Science and Technology (DST), Government of India. BM would like to also acknowledge the support received from the JC Bose Fellowship of the Anusandhan National Research Foundation (ANRF), Government of India. We also acknowledge the use of the Garuda and Kannad HPC cluster facilities at School of Physical Sciences, NISER.
Appendix A Validation of reactor CENS differential rate
We validated our CENS differential event rate modeling for the KKL reactor (Switzerland) using the experimentally observed reactor antineutrino spectrum from the CONUS+ experiment. CONUS+ is a dedicated CENS search experiment employing kg-scale ultra-low-threshold (160-180 eVee) HPGe detectors to measure neutrino–nucleus interactions. The setup is located approximately 20.7 m from the KKL reactor core and has reported the first observation of reactor antineutrino-induced CENS signals with a significance of 3.7 CONUS+_first_observation_2025.
To cross-check our reactor CENS model, we computed the expected differential event rate for a germanium detector placed at 20.7 m from the KKL core and compared it with the CONUS+ measured spectrum. Since the HPGe detectors in CONUS+ measure ionization signals, the nuclear recoil energy from CENS interactions undergoes quenching. Accordingly, we converted the predicted nuclear recoil spectrum into the corresponding electron-equivalent energy using the Lindhard model with .
As shown in Fig. 10, our predicted differential event rate exhibits good agreement with the CONUS+ experimental results, thereby validating both our reactor antineutrino flux model and the CENS differential spectrum modeling in terms of electron-equivalent measurable energy.
Appendix B Comparison of baseline neutrino floor
We compared our baseline neutrino floor results—i.e., the neutrino floor without the effect of reactor antineutrinos—with previously published results for a Ge detector. The comparison was performed for both the discovery-limit–based Billard:APPEC_report and opacity-based OHare:2021utq neutrino floors and shown in Fig. 11.
For the discovery-limit case, our result is consistent with the APPEC Committee Report (2021) Billard:APPEC_report, except near a WIMP mass of . This deviation originates from the significantly reduced systematic uncertainty used for the 8B solar-neutrino flux in our study (2%, see Table 1) compared to the 16% used in the APPEC report. As discussed in Fig. 5, a reduction in systematic uncertainty by an order of magnitude can lower the neutrino floor, thereby expanding the discovery potential for direct dark matter detection. We have verified that adopting a 16% uncertainty for the 8B flux raises the floor, reproducing the result reported by the APPEC community.
For the opacity-based neutrino floor, our baseline result shows reasonable agreement with the published calculation of Ref. OHare:2021utq for a Ge detector. It is worth noting that the O’Hare calculation was performed for the Gran Sasso site, accounting for both geo- and reactor antineutrino contributions at that site. The reactor antineutrino component was found to be negligible owing to the large distances to nearby reactors (). This is consistent with our findings, which indicate that such distant reactors have an insignificant effect on the neutrino floor.